### v1

# Remote Auditing: Design-Based Tests of Randomization, Selection, and Missingness with Broadly Accessible Satellite Imagery

Abstract
Randomized controlled trials (RCTs) are the benchmark for causal inference, yet field implementation can deviate. We here present a remote audit—a design-based, preregistrable diagnostic that uses only pre-treatment satellite imagery to test whether assignment is independent of local conditions. The conditional randomization test of the remote audit evaluates whether treatment assignment is more predictable from pre-treatment satellite features than expected under the experiment’s registered mechanism, providing a finite-sample valid, design-based diagnostic that requires no parametric assumptions. The procedure is finite-sample valid, honors blocks and clusters, and controls multiplicity across image models and resolutions via a max-statistic. We illustrate with two RCTs: Uganda’s Youth Opportunities Program, where the audit corroborates randomization and flags selection and missing-data risks; and a school-based trial in Bangladesh, where assignment is highly predictable from pre-treatment features relative to the stated design, consistent with independent concerns about irregularities. Remote audits complement balance tests, lower early-stage costs, and enable rapid design checks when baseline surveys are expensive or infeasible.
Keywords: Randomization integrity; Satellite imagery; Conditional randomization test; Field experiments
Introduction
“Nullius in verba.”—Motto of the Royal Society (from Horace, Epistles I.i)
Randomized experiments have transformed empirical social science by offering credible causal leverage in hard-to-study environments (Rubin, 2005). In practice, however, the path from a pre-registered randomization mechanism to realized treatment assignment is sometimes fraught. Geography, logistics, bureaucratic discretion, and political pressures can all perturb assignment away from the intended design (Glennerster and Takavarasha, 2013; Olken, 2015). Even modest deviations can matter for inference, particularly when assignment correlates with contextual features that also shape outcomes (Battisti, 2017). Traditional safeguards such as centralized randomization draws, sealed lists, enumerator training, together with ex post diagnostics on covariate balance, manipulation checks, are indispensable (Morgan and Rubin, 2012; Bruhn and McKenzie, 2009). Yet they can be expensive, delayed, or underpowered in the very settings where field experiments are most valuable: low-resource environments, multi-jurisdiction programs, or government deployments in which baseline surveys are difficult or expensive to field at scale (Bouguen et al., 2019).
We here propose a complementary tool: a remote audit of randomization integrity that relies solely on pre-treatment satellite imagery. The idea is simple. If treatment assignment is truly as-if random under the registered mechanism, then it should not be predictable from covariates extracted from images collected prior to randomization. Conversely, if implementers implicitly targeted more accessible, wealthier, less conflict-prone, or otherwise distinctive places—attributes that often leave visual traces at even moderate resolutions—then a predictive signal should be detectable in the imagery. Because satellite archives are global, historical, and increasingly easy to process, this diagnostic can be run early and cheaply, in parallel with—or well in advance of—baseline data collection.
Our approach operationalizes this idea as a conditional randomization test (CRT) (Candes et al., 2018; Hennessy et al., 2016) tailored to field experiments. We (i) extract features from strictly pre-treatment images (e.g., Landsat, Sentinel) using interpretable indices (e.g., nightlight) or off-the-shelf backbones (e.g., CLIP-like encoders, ViT, Swin), (ii) train a predictive model of treatment using only these pre-treatment embeddings, summarize fit with an out-of-sample log-likelihood improvement statistic, and (iii) compare the observed statistic to its finite-sample reference distribution obtained by resampling from the known randomization scheme (honoring blocks, clusters, and treatment fractions). We further provide a max-statistic procedure (Westfall and Young, 1993) to adjust inference across multiple image models, resolutions, and patch sizes, and we discuss simple alternatives like Bonferroni or BH-style control (Thissen, Steinberg and Kuang, 2002).
The remote audit is design-based: validity does not hinge on outcome models or parametric assumptions, only on resampling from the registered assignment mechanism. It uses no post-treatment variables, thereby avoiding “bad control” pitfalls (Angrist and Pischke, 2009; Pearl, 2009; Cinelli and Hazlett, 2020), and because we never condition on image features in the outcome model, we avoid mediator/collider concerns that arise if images are measured after treatment. The audit complements standard balance tests: whereas balance checks examine low-dimensional covariates (often unavailable ex ante), the audit leverages high-dimensional, pre-existing visual context that is ubiquitous and pre-treatment.
We illustrate with two audits. First, re-analyzing Uganda’s government-run Youth Opportunities Program (YOP) RCT (Blattman, Fiala and Martinez, 2014) using only pre-2008 imagery, we find the observed assignment is no more predictable than resamples under the reported lottery, consistent with proper randomization. The same workflow highlights (i) strong predictability of trial participation relative to a national frame and (ii) image-predictive missingness, flagging external validity and data-loss risks. Second, for a school-based RCT in Bangladesh (Begum, Grossman and Islam, 2022), cluster assignment of treatment is itself highly predictable from pre-treatment features relative to the reported design—evidence consistent with independent concerns about irregularities (Bonander et al., 2025). These low-cost audits can shape fieldwork priorities, measurement strategies, and pre-analysis plans.
Our contribution is fourfold. First, we formalize a preregistrable CRT for randomization integrity explicitly adapted to geospatial experiments and to off-the-shelf computer vision embeddings. Second, we provide practical guidance on implementation: pre-treatment image selection, patching and scale, sample-splitting to prevent overfitting, multiple-testing control across models/resolutions, and transparent reporting. Third, we show empirically that remote audits can be informative where baseline covariates are unavailable, delayed, or incomplete—a common situation in development and governance experiments (Dreher and Lohmann, 2015; BenYishay, DiLorenzo and Dolan, 2022; Weisberg, 2009). Fourth, we release a free and coding-free audit online computer application that experimenters can use to perform audits at scale (URL: https://huggingface.co/spaces/cjerzak/ra).
In brief, remote audits will not detect all implementation problems. Many political and social processes are not visible from space, and clouds, revisit cycles, and spatial resolution impose constraints. Nonetheless, as a minimally invasive, design-based check, a remote audit can serve as an early warning system, flagging potential risks and guiding remedial steps before costly downstream investments.
1 A Conditional Randomization Test for Remote Audits
1.1 Design Setup
Let $\Omega$ denote the registered randomization procedure (e.g., complete randomization at rate $\bar{a}$, or stratified randomization within blocks with fixed treatment counts). Consider experimental units $i=1,\ldots,n$ (e.g., villages, neighborhoods, clinics), each with geospatial coordinates $\ell_{i}\in\mathbb{R}^{2}$. Prior to any intervention, we extract a pre-treatment image patch of size $s>0$ centered on $\ell_{i}$ from a public image archive (e.g., Landsat), denoted $\mathbf{M}_{i}=f_{M}(\ell_{i},s)$. From $\mathbf{M}_{i}$, we compute a representation, $\boldsymbol{\phi}_{i}=f_{\boldsymbol{\phi}}(\mathbf{M}_{i})\in\mathbb{R}^{d}$ using either interpretable indices (e.g., vegetation or texture), task-agnostic image embedding models (e.g., CLIP-like encoders, ViT, Swin), or both. [^1] Embeddings are generated variables. When used as regressors for causal estimation they can require additional care (Battaglia et al., 2024). Our design-based test avoids such complications because we only use embeddings to form a test statistic for assignment, not to adjust outcome models. Let $A_{i}\in\{0,1\}$ denote treatment assignment and $Y_{i}$ the outcome.
Under the experiment’s intended design, assignment is independent of all pre-treatment variables, including $\boldsymbol{\phi}_{i}$:
$A_{i}\;\perp\!\!\!\perp\;\boldsymbol{\phi}_{i}$
conditional on the known randomization scheme $\Omega$ (e.g., complete randomization at rate $\bar{a}$, or stratified/block randomization within pre-defined strata). Informally, $\boldsymbol{\phi}_{i}$ should not help predict $A_{i}$ beyond the treatment fraction implied by $\Omega$. If $\boldsymbol{\phi}_{i}$ does predict $A_{i}$ in the observed data substantially better than it typically does under draws from $\Omega$, that is evidence that the realized assignment is atypical of the design—consistent with implementer discretion, operational constraints, or administrative errors aligning treatment with visual correlates of local conditions.
In contrast to covariate balance tests, which examine a low-dimensional set of pre-specified variables, remote audits leverage high-dimensional pre-treatment information available almost everywhere on Earth. [^2] Earth-observation missions provide global coverage and long archives, though revisit, cloud cover, and licensing constraints matter (Barnum, 2022; Tao et al., 2016; Townsend, 2021). Because $\mathbf{M}_{i}$ is measured before treatment on a daily basis for all geographical areas, we avoid post-treatment bias (Angrist and Pischke, 2009; Pearl, 2009); because we never condition on $\boldsymbol{\phi}_{i}$ in outcome analysis, we avoid introducing bad controls or colliders.
[…]
Figure 1: Remote audit intuition. Is latent context $U_{i}$, proxied by imagery, correlated with treatment assignment? The dark, thicker arrow marks the edge whose (effective) presence would make assignment $A_{i}$ predictable from pre-treatment imagery via embeddings, i.e., the path $\boldsymbol{\phi}_{i}\!\to\!A_{i}$. Under the registered mechanism, $A_{i}\perp\!\!\!\perp\boldsymbol{\phi}_{i}$.
1.2 Test Statistic
To turn this intuition into a test, we need a single, learner-agnostic statistic that captures how much the pre-treatment image information predicts treatment assignment beyond what the registered mechanism $\Omega$ would yield by chance. The desiderata are simple: evaluate strictly out of sample to prevent overfitting; anchor the scale to $\Omega$ so that “no signal” maps asymptotically to zero; allow additivity across units and folds for transparent cross-fitting; and avoid dependence on any particular classifier. A natural choice satisfying these criteria is the improvement in the predictive log-likelihood of $A$ on held-out data relative to the baseline assignment probabilities implied by $\Omega$ (or, in stratified/clustered designs, the unit-level baselines $q_{i}$). We treat the embedding–learner pair as a black box that outputs predicted probabilities and summarize only their incremental information. We now define this statistic formally; the Appendix derives its finite-sample reference distribution under $\Omega$.
First, split the sample at random into train and test folds (or use $K$-fold cross-fitting). Using only the training fold, fit any predictive model of $A_{i}$ on $\boldsymbol{\phi}_{i}$ (e.g., ridge-penalized logistic regression, gradient-boosted trees, or a neural model). Denote the probability predictions on the test fold by $\hat{\pi}_{i}=\widehat{\Pr}(A_{i}=1\mid\boldsymbol{\phi}_{i})$. Define the out-of-sample log-likelihood:
$\mathcal{L}\;=\;\sum_{i\in\text{test}}\left(A_{i}\log\hat{\pi}_{i}+(1-A_{i})\log(1-\hat{\pi}_{i})\right).$
To facilitate interpretation, subtract the baseline log-likelihood under the marginal treatment probability $\bar{a}$ implied by $\Omega$:
$T\;=\;\mathcal{L}\;-\;\sum_{i\in\text{test}}\big(A_{i}\log\bar{a}+(1-A_{i})\log(1-\bar{a})\big).$ (1)
Under genuine randomization, we expect $T$ to approach zero (from below) as the learner predicts the baseline treatment assignment with more precision (so that $T\to\mathcal{L}-\mathcal{L}=0$). Substantially positive values indicate that $\boldsymbol{\phi}_{i}$ predicts $A_{i}$ better than chance.
In more complex designs, treatment assignments may vary for each unit (e.g., within blocks), the baseline log-likelihood improvement can be rewritten using unit baseline assignment probabilities $q_{i}\;\equiv\;\Pr_{A\sim\Omega}(A_{i}=1)$:
$T\;=\;\sum_{i\in\text{test}}\Big[A_{i}\log\hat{\pi}_{i}+(1-A_{i})\log(1-\hat{\pi}_{i})\Big]\;-\;\sum_{i\in\text{test}}\Big[A_{i}\log q_{i}+(1-A_{i})\log(1-q_{i})\Big],$ (2)
The log-likelihood improvements in Eq. 1 and Eq. 2 are instances of strictly proper scoring rules (Gneiting and Raftery, 2007), ensuring that increases in $T$ correspond to genuine predictive gains on the test fold (which are assessed against the randomization reference). Alternative proper scores (e.g., Brier improvement, defined as the mean squared error of the predicted probability versus the outcome) are admissible; we use the likelihood scale here because it accumulates naturally across folds and aligns with design-based resampling.
1.3 Reference Distribution and $p$-Values
To interpret the statistic $T$, we need a reference distribution that encodes the experiment’s null: under the registered mechanism $\Omega$, assignment is independent of all pre-treatment information.
The cleanest benchmark is randomization-based. We redraw assignment vectors from $\Omega$ while holding fixed everything that is genuinely pre-treatment—the imagery, the derived embeddings, all preprocessing choices, and the train/test split used for cross-fitting—and for each redraw we refit the learner on the training fold and evaluate on the test fold to recompute the same likelihood-improvement statistic. Because the resampling respects block and cluster constraints by construction, the resulting reference distribution is calibrated in finite samples and requires no parametric approximation. Most importantly, the realized statistic and its resampled counterparts are exchangeable, so the rank of the observed $T$ among them yields a valid measure of extremeness (a $p$-value). This is the essence of the conditional randomization test: a design-based calibration that turns high-dimensional predictability into evidence about departures from the intended randomization (Candes et al., 2018; Hennessy et al., 2016). We now make this construction explicit.
For $b=1,\dots,B$, resample treatment vectors $A^{(b)}\sim\Omega$ subject to the same constraints (block sizes, treatment quotas), re-fit the predictive model on the training fold with $(\boldsymbol{\phi},A^{(b)})$, re-compute $\hat{\pi}^{(b)}$ on the test fold, and compute $T^{(b)}$ exactly as above. The CRT $p$-value is
$p\;=\;\frac{1}{B+1}\left(1+\sum_{b=1}^{B}\mathbf{1}\{T^{(b)}\geq T\}\right),$
which is valid in finite samples for arbitrary test statistics provided the resampling respects $\Omega$ (Candes et al., 2018; Hennessy et al., 2016). This procedure is nonparametric (no asymptotics needed), transparently honors blocking/stratification, and accommodates any off-the-shelf learner as the statistic’s engine. See Appendix for a simple sketch of validity.
An important implementation detail concerns re-fitting versus re-using the trained learner. For strict finite-sample validity, we must refit within each resample using the same cross-fitting protocol (assuming the learner depends on the labels). In moderate samples with simple learners, the computational burden is modest.
Because the remote audit compresses pre-treatment imagery into high-dimensional embeddings $\boldsymbol{\phi}_{i}\in\mathbb{R}^{d}$ (with $d$ often in the hundreds or thousands, arising from multispectral indices and off-the-shelf encoders), the primary computational burden is re-fitting learners within each resample while preserving the design-based calibration to $\Omega$. We can therefore precompute $\boldsymbol{\phi}$ once and apply variance-reduction devices that do not alter the reference distribution.
The first variance reduction device is “common random numbers” (Wright and Ramsay Jr, 1979): fix the cross-fitting split $S$ and all model-training seeds across resamples so that variation in $T^{(b)}$ arises solely from the assignment draws $A^{(b)}\sim\Omega$. This keeps the statistic exchangeable with its resampled counterparts while achieving target Monte Carlo precision with far fewer iterations. Second, in designs that treat exactly half of the units within each block, pairing every draw with its blockwise complement implements an antithetic coupling that further reduces Monte Carlo variance, provided the complement map preserves $\Omega$. Finally, parallelization across $b$ is straightforward. These choices keep the high-dimensional, imagery-driven audit inexpensive and fast without diluting its guarantees: $T$ continues to measure how much the embeddings predict assignment, and its extremeness is still judged against the finite-sample distribution implied by the registered mechanism.
What about multiple testing in the remote audit? If several embeddings or hyperparameters are considered (e.g., CLIP-like encoders, ViT, Swin; patch sizes; resolutions), remote audits should control the family-wise error rate (FWER) or the false discovery rate (FDR). A simple and powerful approach for FWER is the Westfall–Young max-$T$ correction (Westfall and Young, 1993): at each resampled assignment $A^{(b)}$, compute every model’s $T^{(b)}$ and record the maximum. Compare the observed $T$ for each model to this max distribution to obtain adjusted $p$-values. Alternatively, Bonferroni or Benjamini–Hochberg-style methods are available (Thissen, Steinberg and Kuang, 2002). We recommend preregistering the model set and correction rule.
Putting the statistic $T$ and its randomization-based reference under $\Omega$ together, power is likely greatest when deviations from the registered mechanism align with pre-treatment visual signals, when the number of units is larger, when the embedding–learner pair captures stable structure while being evaluated strictly out of sample, and when the design leaves genuine headroom for predictability (highly constrained blockwise quotas leave less to detect). These ingredients translate directly into implementation choices: preserve block totals or cluster labels when resampling under $\Omega$, and prevent spatial leakage in model evaluation by constructing folds that leave out spatial neighborhoods or entire clusters; in our case study we mirror the blocking structure in the folds. Clustered or geographically stratified designs therefore pose no challenge for validity—both the resampling and the cross-fitting simply inherit the design’s constraints.
When assignment depends on factors that satellites cannot see (e.g., patronage networks), the audit will have limited power and this should be reported plainly. By contrast, when assignment comoves with roads, settlement density, roof materials, or land cover—features that proxy accessibility and wealth and are reliably visible from space (Henderson, Storeygard and Weil, 2012; Jean et al., 2016; Watmough et al., 2019; Burke et al., 2021)—the test is often informative even with moderate $n$. In practice, precompute $\boldsymbol{\phi}$ once, start with simple learners and spatially robust cross-fitting, and use the Westfall–Young max-$T$ correction to discipline exploration across representations and scales. A small $p$-value then reads as evidence that the realized assignment is atypical of $\Omega$; a large $p$-value indicates that, at the resolution of the imagery and models used, the audit detected no image-aligned departures from the registered design.
2 A Practical Workflow and Reporting Standards
Having defined the statistic and its design-based calibration, we now turn to practice. The workflow below translates the conditional randomization test into a preregistrable recipe that can be run before, during, or after field mobilization.
Step 1: Define the design.
Record the experiment’s randomization mechanism $\Omega$: complete/stratified/clustered, treatment fractions, and any constraints. If stratified, list strata membership for each unit. These inputs define the resampling.
Step 2: Acquire strictly pre-treatment imagery.
Select images that unambiguously precede any treatment or mobilization. When archives are sparse or cloudy, use compositing or median mosaics across pre-treatment windows. Avoid sensors whose earliest availability is post-treatment. Using post-treatment imagery risks mediator/collider bias if later repurposed in outcome models (Angrist and Pischke, 2009; Pearl, 2009). Archive scene IDs and acquisition dates in the replication package.
Step 3: Construct embeddings.
Choose interpretable indices (e.g., vegetation, built-up proxies) and/or general-purpose backbones (CLIP-like encoders, ViT, Swin) that have been used in Earth observation (Li et al., 2020; Dosovitskiy et al., 2020; Liu et al., 2021). Pretraining on remote-sensing corpora can help but is not required. Document patch size $s$, resolution, bands, and normalizations. Because embeddings are generated variables, keep them fixed throughout the audit and share code for reproducibility (Battaglia et al., 2024).
Step 4: Fit the predictive model with sample-splitting.
Use simple learners first (e.g., tree-based models), then consider neural models. Always evaluate out-of-sample (held-out fold or cross-fitting). Save the likelihood-based statistic $T$.
Step 5: Resample under $\Omega$ and compute the max-$T$.
Draw $B$ assignment vectors consistent with $\Omega$ (e.g., B=1, [^0] B=1{,}000), recompute $T^{(b)}$. If multiple image embedding representations are used, record the maximum test statistic across model per resample. Report adjusted $p$-values and show the observed $T$ against the reference distribution.
Step 6: Interpret cautiously and report transparently.
A small $p$-value suggests an atypical assignment relative to $\Omega$, consistent with implementation deviations; a large $p$-value does not prove correct execution, only that the audit detected no image-aligned deviations. Report model choices, pre-treatment windows, resampling details, and multiple-testing adjustments. Provide code and hashes for imagery products to support reproducibility.
Auxiliary audits.
The same logic supports audits of (i) selection into the experiment (predicting membership among a broader frame) and (ii) missingness (predicting which units have missing variables). These are not design-based in the same sense as the randomization audit because the resampling reference is less well pinned down; we therefore treat them as descriptive early-warning diagnostics that can motivate reweighting, oversampling, or targeted field follow-up.
Item Recommendation
$\square$ Design Describe $\Omega$: complete/stratified/clustered; treatment fractions; any constraints.
$\square$ Pre-treatment window Commit to dates and sensors that strictly precede treatment. Document cloud handling and compositing.
$\square$ Embedding set Pre-specify models (e.g., CLIP-like, ViT, Swin) and interpretable indices; fix patch size(s) $s$.
$\square$ Evaluation Use sample-splitting or cross-fitting; define $T$ as out-of-sample log-likelihood improvement.
$\square$ Resampling Set $B$ (e.g., 1,0001{,}000) and honor blocks/clusters in draws from $\Omega$.
$\square$ Multiplicity Use Westfall–Young max-$T$; alternatively, Bonferroni/BH with justification.
$\square$ Outputs Report adjusted $p$-values, reference distributions, and observed $T$; archive code and imagery.
$\square$ Auxiliary audits If used, label as descriptive diagnostics (selection, missingness) and report separately.
$\square$ Ethics & transparency Avoid post-treatment data; ensure geospatial privacy and responsible disclosure.
Table 1: Checklist for preregistering and reporting a remote audit.
Selection into the study frame.
Let $S_{i}\in\{0,1\}$ indicate whether unit $i$ in a broader, policy-relevant universe is enrolled in the experimental sample. When $\boldsymbol{\phi}_{i}$ strongly predicts $S_{i}$, the enrolled sample differs systematically from the target universe along pre-treatment features that are visible from space, raising an external-validity warning even if within-sample randomization is sound. We operationalize this as a covariate-shift diagnostic: train a classifier to distinguish enrolled units from units drawn from the putative frame using only $\boldsymbol{\phi}$, evaluate strictly out of sample, and summarize fit via likelihood improvement relative to the marginal sampling rate, and permute or randomly re-draw the $S_{i}$ selection indicator. Because there is no registered or otherwise defensible resampling mechanism for $S$ analogous to $\Omega$, these quantities are reported as descriptive diagnostics rather than design-valid tests.
Missingness and data quality.
Let $R_{ij}\in\{0,1\}$ indicate whether variable $j$ is observed for unit $i$. If $\boldsymbol{\phi}_{i}$ predicts $R_{ij}$ out of sample, then complete-case analyses risk bias because missingness correlates with pre-treatment context that may also shape outcomes, enumerator access, or compliance. We therefore fit response models $\hat{\rho}_{ij}(\boldsymbol{\phi}_{i})=\widehat{\Pr}(R_{ij}=1\mid\boldsymbol{\phi}_{i})$ using the same cross-fitting protocol and summarize predictiveness on the likelihood scale relative to the marginal response rate $\bar{r}_{j}$, with descriptive randomization inference. A strong signal can help focus field efforts: high-risk or missingness locations can be prioritized for follow-up, instruments can be adapted for hard-to-reach contexts, and data collection modes can be diversified before surveys are fully fielded. Because imagery is strictly pre-treatment and available daily, these diagnostics can be updated in real time during enumeration without peeking at outcomes.
When analysis requires adjustment, the same response models can be pre-specified as building blocks for principled corrections that do not rely on post-treatment information. For variables where missingness is plausibly at random given $\boldsymbol{\phi}$, inverse-probability weights $w_{ij}=R_{ij}/\hat{\rho}_{ij}(\boldsymbol{\phi}_{i})$ or multiple imputation models that condition on $\boldsymbol{\phi}$ provide transparent remedies; doubly robust procedures that combine a response model with an outcome model can be declared in the pre-analysis plan and implemented without altering the design-based logic of the randomization audit (Seaman and Vansteelandt, 2018). Where missingness is likely non-ignorable even after conditioning on $\boldsymbol{\phi}$, the imagery-based diagnostics still add value by localizing the problem and motivating sensitivity analyses and targeted re-contact. As with selection, we report these quantities as diagnostics rather than as hypothesis tests, and we apply the same cross-validation, sample-splitting, and multiplicity discipline used elsewhere in the audit.
In sum, selection and missingness audits repurpose the remote-audit machinery to answer two practical questions that matter for interpretation and planning: who entered the study relative to the intended universe, and where the data are likely to be thin. Their outputs—likelihood improvements, calibrated risk scores, overlap assessments, and, when warranted, transport or response weights—are easily preregistered, cheap to compute, and actionable during fieldwork. They do not substitute for theory or for direct measurement, and they are not design-valid tests in the sense of the conditional randomization procedure. They are, instead, early-warning indicators that help align designs with their targets and protect inference against predictable patterns of data loss.
3 Case Study: A Remote Audit of the Youth Opportunities Program (Uganda)
We now apply the remote audit to the government-run Youth Opportunities Program (YOP), launched in 2008 in Uganda (Blattman, Fiala and Martinez, 2014). Groups of young adults submitted business plans for cash grants; a lottery determined recipients. The trial has been widely cited and influential. We ask whether pre-treatment satellite imagery—without any survey covariates—could have verified randomization and flagged potential issues regarding selection or data missingness early. In this case, we know of no known reports of randomization problems.
Units and imagery.
We treat applicants’ villages (geocoded from administrative names) as units. We extract pre-2008 image patches from Landsat archives, which are image composites to mitigate clouds and speckle issues. We compute embeddings from an off-the-shelf backbone used in remote sensing using an EO-fine-tuned CLIP model (Li et al., 2020). Patch sizes span the village and immediate environs to capture accessibility and settlement structure.
Figure 2: Experimental frame. Geocoded locations of Youth Opportunities Program (YOP) units in Uganda. The map illustrates nationwide dispersion across settlement types; shaded areas are for orientation only.
Design and resampling.
We reconstruct the reported randomization scheme (e.g., treatment fractions) from the published record (Blattman, Fiala and Martinez, 2014). The CRT resamples treatment vectors consistent with these constraints. For each draw, we recompute the test statistic.
Results.
Figure 3 reports three diagnostics. The randomization audit (Panel A) shows the observed likelihood-improvement statistic (vertical line) within the reference distribution under $\Omega$; the $p$-value is greater than 0.05, consistent with random assignment not being more predictable from pre-treatment imagery than chance. Two auxiliary checks are also informative. The selection audit (Panel B) contrasts YOP villages against a national frame: image features sharply distinguish enrolled vs. unenrolled locations, suggesting external validity concerns if these differences moderate treatment effects. Finally, the missingness audit (Panel C) indicates that pre-treatment features predict which units have missing variables, suggesting non-random data loss that merits attention in analysis plans (e.g., pre-specified handling of attrition).
Panel A: Randomization audit
Panel B: Selection audit
Panel C: Missingness audit
Figure 3: Remote audit results. Each panel displays the reference distribution of the max-statistic obtained from resampling the relevant process (randomization in Panel A; sampling frame or missingness mechanism in Panels B–C) and marks the observed value (vertical line). In Panel A, the observed assignment is not more predictable from imagery than draws from the reported randomization, consistent with integrity of the lottery (Blattman, Fiala and Martinez, 2014). Panels B–C highlight auxiliary risks to external validity and systematic missingness.
Interpretation and use.
In this application, the remote audit would have (i) corroborated the lottery before expensive baseline enumeration, (ii) offered an early warning about representativeness (selection audit), and (iii) prompted pre-analysis plans for handling missing data differentially by location. None of these requires survey covariates. The audit also does not preclude standard balance tests once baselines are collected. Instead, it provides a low-cost, early-stage diagnostic that can be embedded in preregistration and updated as imagery and geocodes are refined.
4 Case Study: Detecting Faulty Randomization Based on a Retracted RCT in Bangladesh
We now apply the remote audit to Begum, Grossman and Islam (2022). Bonander et al. (2025) have raised substantial concerns about the randomization in this setting—including evidence that treatment assignment coincided perfectly with administrative boundaries and that the randomization is identical to some of the study authors used in prior, now retracted work (Islam, 2019). Our audit asks a narrower, design-based question: is the realized assignment in Begum, Grossman and Islam (2022) more predictable from pre-existing local conditions visible from satellite imagery than draws from the study’s purported randomization? The answer is yes, and strongly so, which is consistent with the broader record of irregularities summarized in Bonander et al. (2025). If the answer is yes, then investigators could have deployed the remote audit early on to identify problems with the RCT implementation before incurring the expense of on-the-ground data collection.
Setup.
[…]
Results.
[…]
Interpretation.
The CRT does not identify who or what induced the deviation, nor does it imply that imagery features were used in implementation. It establishes a finite-sample discrepancy: under the claimed design, assignment should not be recoverable from pre-existing landscape signals; yet it is. Coupled with the independent evidence of geographically clustered assignment and shared treatment vectors across linked projects, the audit supports the conclusion that the realized allocation in Begum, Grossman and Islam (2022) was inconsistent with the stated randomization protocol. As with any imagery-only diagnostic, undetected deviations that are orthogonal to what satellites can see remain possible; here, however, two distinct learners point to the same rejection, and the effect sizes are large relative to the null. This analysis shows how aspects of randomization integrity can be tested before any fieldwork begins using the remote audit, successfully detecting possible irregularities noted in the replication community.
[…]
5 What Else Should Analysts Check in Remote Audits?
Design-valid $p$-values are a starting point rather than a stopping rule. Because a remote audit inevitably reflects a handful of investigator choices, it is natural to ask whether the signal persists under reasonable alternatives. The brief checks below are quick to run and help separate substantive evidence from artifacts of modeling or preprocessing.
Vary the statistic and the data split.
Investigators can evaluate alternative strictly proper scores—such as Brier improvement or average cross-entropy per unit—to confirm that conclusions do not hinge on the score’s scale. It is also prudent to examine stability across random seeds and cross-fitting schemes, including spatially robust folds (e.g., leave-one-region-out or leave-one-cluster-out). Reporting the dispersion of $T$ (for example, an interquartile range across repeats) makes the degree of stability transparent.
Probe scale, sensors, and representation.
Choices about spatial scale and sensing are often consequential. Analysts may explore a range of patch sizes $s$ that plausibly cover geocoding error and local spillovers, and, where multiple pre-treatment sensors are available, run parallel analyses for each (with bandpass harmonization as needed). Comparing interpretable indices (vegetation, built-up, texture) with learned encoders helps distinguish model idiosyncrasies from image-aligned signal and can ease communication with non-technical audiences when the two agree.
Placebos and stress tests.
Structure-destroying placebos—such as random rotations of patches or within-region permutations of embeddings—help verify that the pipeline does not manufacture predictability. Complementary stress tests can inject stylized deviations that mimic realistic implementation failures (e.g., favoritism toward road-adjacent or administrative-boundary units) to assess whether the audit would detect such problems at application-relevant sample sizes. A simple simulation calibrated to the study provides a practical power check.
When the audit suggests potential issues, investigators typically consider complementary design or analysis steps:
  • Randomization flagged. Teams often tighten operational constraints or add stratification in a re-randomization, strengthen monitoring of assignment processes (seeds, logs, audit trails), and document the deviation together with pre-specified adaptations for inference (e.g., design-based variance, randomization inference, or sensitivity analysis). If re-randomization occurs, the audit is usually re-run and archived alongside the final assignment.
  • Selection flagged. It can be useful to revisit the sampling frame and clarify the target-population estimand, consider reweighting or targeted enrollment to improve coverage of under-represented areas, and state limits to external validity transparently; transported or reweighted estimates may be offered as secondary analyses where appropriate.
  • Missingness flagged. Common responses include additional follow-up in hard-to-reach locations, small adaptations to instruments or field protocols, and pre-specification of imputation or inverse-probability weighting (avoiding complete-case analyses when missingness is predictably related to $\boldsymbol{\phi}$); attrition bounds provide an additional robustness assessment.
Most of these adjustments are far cheaper and cleaner before enumeration begins. That is the practical advantage of an imagery-only audit: it surfaces potential problems early enough that designs can still be improved rather than merely documented.
6 Scope, Risks, and Implementation Guidance
Remote Audit in Context.
Researchers already exploit satellite imagery to measure outcomes, build covariates, and study heterogeneity (Jean et al., 2016; Yeh et al., 2020; Jerzak, Johansson and Daoud, 2023; Torres and Pugh, 2022). Imagery has also been used model-based to mitigate confounding when pre-treatment signals proxy latent conditions (Sanford, 2021; Burke et al., 2021). Our contribution is orthogonal to these uses. We deploy imagery in a design-based capacity: a conditional randomization test that asks whether realized assignment ($A_{i}$) is independent of pre-treatment embeddings ($\boldsymbol{\phi}_{i}$) under the registered mechanism ($\Omega$). Because the audit never adjusts outcomes, it avoids “bad control” pitfalls and mediator/collider concerns that attend post-treatment measurement in model-based pipelines (Angrist and Pischke, 2009; Pearl, 2009; Cinelli and Hazlett, 2020). Where model-based approaches typically require identification assumptions and sensitivity analysis (Rosenbaum and Rubin, 1983; Oster, 2019), our audit derives validity from the design itself: extremeness is judged against the experiment’s finite-sample reference generated by draws from ($\Omega$). In this sense, the paper reframes what imagery is for—not as a control set inside an outcome model, but as a ubiquitous, pre-existing source of design-stage information capable of certifying whether assignment behaves as if random.
What satellites can test—and what they cannot.
Visual signals reliably register elements of the built and natural environment—settlement structure, roads, roof materials, vegetation, hydrology, and some economic activity such as night lights (Henderson, Storeygard and Weil, 2012; Jean et al., 2016; Watmough et al., 2019). These are precisely the kinds of context that often co-move with accessibility, wealth, and administrative capacity. A rejection by our audit therefore indicates an image-aligned deviation from ($\Omega$); it does not identify mechanisms or actors, which require field investigation. Conversely, a non-rejection is not a certificate of perfection: many forces relevant to assignment (patronage networks, norms, internal procedures) leave weak or no satellite trace at available resolutions. The paper’s insistence on strictly pre-treatment imagery is central to this point. Using images captured after mobilization risks encoding mediators (construction, publicity) or conditioning on colliders (selection into measurement), with familiar causal consequences (Angrist and Pischke, 2009; Pearl, 2009). Our guidance is to commit ex ante to a pre-treatment window, to document sensor choice, compositing, and quality control, and to quantify geocoding uncertainty with patch sizes (s) that are robust to plausible error.
Principles for valid and reproducible audits.
The audit’s guarantees hinge on separating representation from inference and evaluating strictly out of sample. We therefore precompute embeddings—interpretable indices and/or off‑the‑shelf encoders—fixing checkpoints, layers, and normalizations ex ante; we share hashes and manifests so that ($\boldsymbol{\phi}$) can be reproduced byte-for-byte. We fit simple learners first, deploying more complicated models only as needed, and use repeated (K)-fold cross-fitting with seeds recorded. Clustered designs should leave whole clusters out of folds to mitigate spatial leakage. The statistic is defined as out‑of‑sample log‑likelihood improvement relative to the assignment probabilities implied by ($\Omega$), so that “no signal” maps transparently to zero. Because our test is randomization‑based, it naturally honors blocks and clusters during resampling, and its pp‑values are finite‑sample valid without parametric assumptions. These elements make the audit naturally preregistrable: the image window, representation, folds, statistic, and resampling rule can all be locked in advance, limiting the researcher’s degrees of freedom (Bouguen et al., 2019).
Multiplicity, computation, and decision rules.
Exploration across embeddings, resolutions, and patch sizes is often scientifically reasonable but inflates Type I error if not considered in analysis. We therefore recommend Westfall–Young max‑$T$ adjustment as a powerful, design‑compatible correction; simple conservative alternatives (Bonferroni, BH‑style control) are available when appropriate (Westfall and Young, 1993; Thissen, Steinberg and Kuang, 2002). Preregistered decision rules should specify thresholds (e.g., $\alpha$ = 0.05) and remediation steps—re‑randomization, tightening operational constraints, or proceeding with transparent documentation when re‑randomization is infeasible—treating the audit as a design‑stage gate akin to power analysis.
Interpretability, communication, and ethics.
Although the audit’s validity does not depend on interpretability, policy audiences benefit from understanding “what the model saw.” Post‑hoc summaries such as permutation importance or representative patch visualizations can aid communication, provided we separate explanation from inference. Finally, because satellite data may be granular, ethical use requires attention to geospatial privacy and local norms. Sharing processed features and manifests rather than raw tiles, when feasible, reduces re‑identification risk without compromising the design-based logic of the audit.
In sum, the remote audit’s distinctive contribution is to repurpose broadly accessible pre‑treatment imagery for a genuinely design‑based question—did realized assignment conform to ($\Omega$)?—and to do so with a preregistrable, finite‑sample‑valid procedure that scales across stratified and clustered experiments. This positions remote audits as a low‑cost complement to conventional diagnostics: powerful when deviations align with visible context, transparent about their limits when they do not, and deployable early enough to improve designs rather than merely document them.
7 Conclusion
We develop and demonstrate a remote audit of randomization integrity that leverages only pre-treatment satellite imagery and a conditional randomization test. The audit is valid in finite samples, easily preregistered, and compatible with stratified and clustered designs. In Uganda’s Youth Opportunities Program (Blattman, Fiala and Martinez, 2014), it would have corroborated the reported lottery while flagging selection and missing-data risks that matter for interpretation and design; in a field experiment in Bangladesh, randomization integrity is itself questioned, in line with concerns raised by an independent team of investigators.
Remote audits are not a panacea. The remote audit detects only those design-context correlations that are captured by pre-treatment imagery. Some factors that influence assignment may be invisible from space or imperfectly proxied at available resolutions. Geocoding errors and cloud cover can reduce power. A non-rejection is not a proof of integrity, and a rejection does not uniquely identify the source of deviation. For selection and missingness extensions, the tests are descriptive rather than design-based. Despite these limits, the audit adds a transparent, low-cost, preregistrable safeguard where ground measurement is difficult.
As global image archives deepen and off-the-shelf vision models improve (Li et al., 2020; Dosovitskiy et al., 2020; Liu et al., 2021), the cost of running such audits is becoming vanishingly small. For experiments in challenging environments—where baseline enumeration is costly, schedules tight, and logistics uncertain—remote audits offer a practical early-warning system. We recommend incorporating them into preregistration and routine experimental quality assurance alongside conventional balance checks and process documentation (Bouguen et al., 2019). $\square$
Appendix
Proposition (Finite-sample validity of the remote audit). Fix the pre-treatment embeddings $\boldsymbol{\phi}=\{\boldsymbol{\phi}_{i}\}_{i=1}^{n}$ and a fold-splitting scheme $H$ (which may be a deterministic function of $\boldsymbol{\phi}$ or drawn independently of $A$). Let $g(\boldsymbol{\phi},A,H)$ be the audit’s statistic—e.g., the out-of-sample log-likelihood improvement $T$ computed by training on the $H$-defined training fold and evaluating on the test fold. Suppose the realized assignment $A$ is drawn from the registered randomization mechanism $\Omega$ and is (by design) independent of all pre-treatment variables, including $\boldsymbol{\phi}$. For $b=1,\ldots,B$, draw $A^{(b)}\sim\Omega$ (independently of each other and of $A$), and define $T=g(\boldsymbol{\phi},A,H)$ and $T^{(b)}=g(\boldsymbol{\phi},A^{(b)},H)$, recomputing the learner under each $A^{(b)}$. Then the $p$-value
$p=\frac{1}{B+1}\Big(1+\sum_{b=1}^{B}\mathbf{1}\{T^{(b)}\geq T\}\Big)$
satisfies $\Pr\!\big(p\leq\alpha\mid\boldsymbol{\phi},H\big)\leq\alpha$ for all $\alpha\in[0,1]$. Hence the remote audit controls Type I error at level $\alpha$ in finite samples, conditional on $(\boldsymbol{\phi},H)$. The result continues to hold for stratified or clustered designs provided $\Omega$ and the resampling preserve the design’s block/cluster constraints.
Proof. Condition on $(\boldsymbol{\phi},H)$. Under the null, $A\sim\Omega$ and $A^{(1)},\ldots,A^{(B)}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\Omega$ are exchangeable. Applying the fixed, measurable map $g(\cdot)$ to each assignment yields exchangeable statistics $\big(T,T^{(1)},\ldots,T^{(B)}\big)$. Therefore the rank of $T$ among these B+1B{+}1 values is uniformly distributed on $\{1,\ldots,B{+}1\}$ (with ties handled by the $\geq$ rule), which implies that $p$ is super-uniform and $\Pr(p\leq\alpha\mid\boldsymbol{\phi},H)\leq\alpha$. When $\Omega$ imposes block or cluster totals, exchangeability holds conditional on those totals, so the same argument applies. $\square$
Remark. Refitting the predictive model within each resample is what ensures that $g(\boldsymbol{\phi},\cdot,S)$ treats every draw from $\Omega$ symmetrically; reusing a learner trained only on the realized AA can break exchangeability. See Candes et al. (2018) and Hennessy et al. (2016) for more information.